OpenStudy (anonymous):

Richard has four times as marbles as john. If Richard gave John 18 marbles they would have the same amount. How many marbles has each?

OpenStudy (anonymous):

let john(j)+richard(r)=x in first case, j:r=1:4 j/r=1/4 4j=r --------------1 in second case, j+18:r-18=1:1 j+18/r-18=1/1 j+18=r-18 j+36=r-------------2 substitute 1 into 2 j+36=4j 36=3j j=12 from 1 4j=r 4(12)=r r=48 so john has 12 marbles; while richard has 48 marbles

OpenStudy (anonymous):

How does j/r become 1/4??

OpenStudy (anonymous):

ratio (:) means divide

OpenStudy (anonymous):

use ratio for long question so that u dont confuse which 1 is bigger ( like Richard has four times as marbles as john) so u may confuse is john bigger or richard... use : so that u know richard is bigger than john

OpenStudy (anonymous):

RICHARD has FOUR times as marbles as john so RICHARD:john=FOUR:one

OpenStudy (anonymous):

got to go.. bye...

OpenStudy (anonymous):

does ratio help with word problems?

OpenStudy (anonymous):

oh ok thanks bye. :)